40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 16 23:25:07 2017, gautam, Update, SUS, MC1 SUS electronics investigation 20170116_231625.pngIMG_7175.JPGIMG_7174.JPG
    Reply  Tue Jan 17 21:29:52 2017, gautam, Update, SUS, MC1 SUS electronics investigation 
       Reply  Wed Jan 18 11:40:54 2017, gautam, Update, SUS, MC1 SUS electronics investigation MC1glitches.pngMC1coilGlitches.pngMC1CoilGlitchZoomed.png
          Reply  Wed Jan 18 14:23:47 2017, gautam, Update, SUS, MC1 SUS electronics investigation 
             Reply  Wed Jan 18 18:44:53 2017, gautam, Update, SUS, MC1 SUS electronics investigation MCsatBoxSwitch.png
                Reply  Thu Jan 19 08:25:12 2017, Steve, Update, SUS, MC1 SUS electronics investigation MC1_MC3_ITMY_ETMX_sensors.pngsensors_UL.png
                   Reply  Thu Jan 19 12:00:10 2017, gautam, Update, SUS, MC1 SUS electronics investigation SatBoxSwitchGlitch.png
                      Reply  Thu Jan 19 19:56:09 2017, rana, Update, SUS, MC1 SUS electronics investigation 
                         Reply  Fri Jan 20 11:16:30 2017, gautam, Update, SUS, MC1 SUS electronics investigation MCglitchesReturn.png
                            Reply  Mon Jan 23 10:24:01 2017, Steve, Update, SUS, MC1 SUS electronics investigation 3glitchingSUS.png
Message ID: 12731     Entry time: Wed Jan 18 11:40:54 2017     In reply to: 12728     Reply to this: 12734
Author: gautam 
Type: Update 
Category: SUS 
Subject: MC1 SUS electronics investigation 

After the repair of the faulty LEMO cable, I left MC1 with it's watchdog off overnight. Unfortunately, it looks like the problem still persists. The first attachment shows a second trend plot for the past 15 hours. Towards the left end of the plot, you can see where I re-connected the LEMO cable for the LR/UR channels.

A couple of months ago, I added a BLRMS block for the IMC optics that calculates BLRMS for the shadow sensor output as well as the coil output. Looking at this trend overnight, I noticed that the glitches appear in the coil outputs as well, as shown in the plot below, which is for a 1 hour stretch last night (I used the full data from a 16Hz coil output channel and not the BLRMS, I am not sure if there is a DQ'ed version of the coil outputs?).

Zooming in further to one of these glitches, we can see that the glitches in the coil and shadow sensor signals are in fact coincident.

But given that the watchdog was turned off all this time, the only voltage going to the coils should be the DC bias voltages. So does this not support the hypothesis that the problem lies in the part of the signal chain that supplies the bias voltage to the coils?

Never mind, the "coil output" channel isn't a true readback of the voltage to the coil, but is the calculated damping output (which is not sent to the coils when the watchdog is shutdown...

ELOG V3.1.3-